complete solutions

(3.33<-145°)*(6.79<97°)

(5e5.6i)/(9.75e-0.37i)



1
Expert's answer
2022-02-14T10:18:38-0500

i)

"3.33e^{i(-145\degree)}\cdot 6.79e^{i(97\degree)}=22.6107e^{i(-48\degree)}"



ii)


"\dfrac{5e^{5.6i}}{9.75e^{-0.37i}}=\dfrac{20}{39}\cdot e^{5.97i}"